Suppose the demand for money is L=0.20Y, the money supply is 200, consumption: C=90+0.80YD, taxes T=50, Investment: I=140-5r, and government purchases: G=50.
i) ii) iii)
Derive the IS and LM equations
Find equilibrium output, and the rate of interest
Estimate the investment level if government spending increases by 20?
1
Expert's answer
2020-08-18T12:54:26-0400

i) IS equation


=Y=C+I+G


=90+0.8(Y-50)+140-5r+50


=90+0.8Y-32+140-5r+50


0.2Y=248-5r


Y=1240-25r


LM equation


First equate the Money Supply with Money Demand;


200=0.20 Y


Y=1,000


ii) The equilibrium level of output Y, =1,000


Rate of interest ;1000=1240-25r


-25r=-240


r=9.6%


iii)Increase in Government spending by 20%=1.2"\times"50=60


1000=850+60+I


I=1000-(850+60)


Investment =90